OpenStudy (anonymous):

Stephanie is organizing the seating arrangements for a concert in her local park. The first row's seats are modeled by the function f(x) = 15. To give everyone a good view, each row must have four more seats than the row before it. The number of additional seats a row has is modeled by the function s(x) = 4(x - 1). Explain to Stephanie how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the fortieth row. Use complete sentences.

OpenStudy (anonymous):

\[F \left( x \right)=f \left( x \right)+s \left( x \right)=15+4\left( x-1 \right),x \in N\]

OpenStudy (anonymous):

could you explain how you got that for me?

OpenStudy (anonymous):

if x=1 F(x)=15+4(1-1)=15+4*0=15 x=2 F(x)=15+4(2-1)=15+4=19 ..................... it is given in first row=15 additional seats in each row=4(x-1)

OpenStudy (anonymous):

ok thanks for the help :D

OpenStudy (anonymous):

yw